OpenStudy (anonymous):

Evaluate. log3 Choose one answer. a. 27 b. 4 c. -27 d. -4

jimthompson5910 (jim_thompson5910):

just log3?

jimthompson5910 (jim_thompson5910):

or is there something more?

OpenStudy (anonymous):

there is 1/81 too

jimthompson5910 (jim_thompson5910):

so the full problem is \[\Large \log_{3}\left(\frac{1}{81}\right)\]

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

d

jimthompson5910 (jim_thompson5910):

\[\Large \log_{3}\left(\frac{1}{81}\right)\] \[\Large \log_{3}\left(\frac{1}{3^4}\right)\] \[\Large \log_{3}\left(3^{-4}\right)\] \[\Large -4\log_{3}\left(3\right)\] \[\Large -4\left(1\right)\] \[\Large -4\] So \[\Large \log_{3}\left(\frac{1}{81}\right)=-4\] Which makes the answer choice D

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

if there is one that says log x=4 how would you solve that

jimthompson5910 (jim_thompson5910):

If it doesn't have a base shown, then it's by default, base 10

OpenStudy (anonymous):

would it just be 4 the options are a. x = 4 b. x = 40 c. x = 1,000 d. x = 10,000

jimthompson5910 (jim_thompson5910):

\[\Large \log(x) = 4\] \[\Large \log_{10}(x) = 4\] \[\Large x = 10^{4}\] \[\Large x = 10,000\] Note: in the third step, I'm converting to exponential form using the idea that \[\Large \log_{b}(x) = y\] converts to \[\Large x = b^y\]

OpenStudy (anonymous):

so the answer is 10,000

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks!

jimthompson5910 (jim_thompson5910):

yw